# Assessing Omitted Variable Bias

### v5

@@ 5.3 Assessing Selection on Observables
Sensitivity Analysis Based on $\delta_{\text{resid}}$
We give a brief summary of the sensitivity analysis based on Oster (2019) (hereafter Oster), as used in Bazzi et al. (2020). They use sensitivity parameter $\delta_{\text{resid}}$, which is informally described in Oster’s appendix A (see appendix C of our companion paper Diegert et al. (2025) for a detailed, formal description of this parameter). The variables $(X,W_{1},W_{2})$ are replaced by $(X^{\perp W_{0}},W_{1}^{\perp W_{0}},W_{2}^{\perp W_{0}})$ to adjust for the non-calibration covariates $W_{0}$. Finally, the auxiliary sensitivity parameter in Oster’s analysis, $R_{Y\sim X,W_{1},W_{2}\mathrel{{\bullet}}W_{0}}^{2}$, is set to $1.3\cdot\widehat{R}_{Y\sim X,W_{1}\mathrel{{\bullet}}W_{0}}^{2}$. Using these parameters and Oster’s analysis, they report the estimated breakdown point, which is a measure of how large $\delta_{\text{resid}}$ must be for the estimated effect to be fully explained away by the unobservables.
The second row of Panel B of table 5.2 shows sample analog estimates of this breakdown point, which is commonly referred to as Oster’s delta. The first row of Panel B shows the values of Oster’s delta as reported in table 2 of Bazzi et al. (2020). These were computed incorrectly. [^4] There are three differences between the two rows in Panel B. First, it appears to us that, rather than using the correct expression in Proposition 3 of Oster (2019), Bazzi et al. (2020) set the first displayed equation on page 193 of that paper equal to zero and solved for $\delta_{\text{resid}}$. That does not give the correct breakdown point. This mistake was not unique to those authors; while replicating other results, we discovered the same mistake in several other papers published in ‘top 5’ economics journals. Second, Bazzi et al. (2020) used $\widehat{R}_{Y\sim X,W_{0},W_{1}}^{2}$ and $\widehat{R}_{Y\sim X,W_{0}}^{2}$ to adjust for $W_{0}$ rather than $\widehat{R}_{Y\sim X,W_{1}\mathrel{{\bullet}}W_{0}}^{2}$ and $\widehat{R}_{Y\sim X\mathrel{{\bullet}}W_{0}}^{2}$. Third, as we discussed earlier, they include the individual demographics and survey wave fixed effects in $W_{1}$ rather than $W_{0}$. This is a valid choice, but we chose to instead include these variables in $W_{0}$ to keep the covariates in $W_{1}$ constant across columns (1)–(5) for comparability. Based on these estimates, Bazzi et al. (2020) conclude:
@@
The estimates in table 2 show a substantial range of correlation between the observed covariates in $W_{1}$. Recall that the exogenous controls assumption says that each element of $W_{1}$ is uncorrelated with $W_{2}$, after partialling out $W_{0}$. Thus if $W_{2}$ was included in this table it would have a value of zero. Therefore, if $W_{2}$ is a variable similar to the components of $W_{1}$ then we would expect exogenous controls to fail. This suggests that, in this application, robustness conclusions based on $\delta_{\text{resid}}$ could be incorrect. Consequently, one should instead use sensitivity analysis methods which are valid even when the controls are endogenous.
@@
Sensitivity Analysis Based on $r_{X}$ and $r_{Y}$
@@
Prefers Cut Public Spending on Poor Republican Presidential Vote Share
@@
Calibrating $\bar{r}_{X}$. Table 5.3 shows estimates of the calibration parameters $\rho_{k}$ that we discussed in section 3.4. Although the treatment variable is the same across all specifications in Table 5.2, the estimation datasets vary. Consequently, the estimates of $\rho_{k}$ vary across specifications as well. Here we only show the estimates corresponding to the specifications in columns (1) and (5) for brevity. First consider column (5), the results for average Republican vote share. Here we see that $\widehat{\rho}_{k}$ is smaller than the $\bar{r}_{X}^{\text{bp}}$ breakdown point estimate of 80.4% for all but one covariate value, suggesting that the results for average Republican vote share is quite robust to selection on unobservables. In contrast, consider column (1), the results for the cut spending on the poor outcome variable. In this case, the $\bar{r}_{X}^{\text{bp}}$ breakdown point estimate of 2.81% is smaller than all the calibration parameters $\widehat{\rho}_{k}$. Even if we use the less conservative breakdown point $\widehat{\bar{r}}^{\text{bp}}=54.1$%, this value is still smaller than half of the calibration parameters. This suggests that the results for cut spending on the poor are sensitive to selection on unobservables. Overall, our calibration analysis confirms the findings discussed above: The conclusions using the behavior based outcomes are substantially more robust than the conclusions using questionnaire based outcomes.
Figure 1: Sensitivity Analysis for Average Republican Vote Share. See body text for discussion.
@@
Next we consider robustness to choices of $\bar{r}_{Y}\neq\bar{r}_{X}$. To show this, we plot the breakdown frontier $\bar{r}_{Y}^{\text{bf}}(\bar{r}_{X},0)$ that we showed how to compute in Theorem 3. This is shown as the solid black line in the right plot of figure 1. For any value of $(\bar{r}_{X},\bar{r}_{Y})$ below this line we can conclude that $\beta_{\text{long}}>0$. For example, $(\bar{r}_{X},\bar{r}_{Y})=(2,0.5)$ lies below this line. This means that our conclusion that $\beta_{\text{long}}>0$ is robust to unobservables up to twice as strong as the observables for their impact on treatment, and up to half as strong as the observables for their impact on outcomes.
In section 3.5 we also extended our results to allow researchers to restrict the magnitude of control endogeneity as well. These restrictions shift the breakdown frontier away from zero, as shown by the gray lines in the right plot of figure 1. These plots correspond to the restriction $R_{W_{2}\sim W_{1}\mathrel{{\bullet}}W_{0}}\in[0,\bar{c}]$ for the various displayed values of $\bar{c}$. For example, if we set $\bar{c}=0.7$ then the point $\bar{r}_{X}=\bar{r}_{Y}=110\%$ is now below the breakdown frontier. Hence with this restriction on control endogeneity, our conclusion $\beta_{\text{long}}>0$ is robust to unobservables up to 1.1 times as strong as unobservables in for their impact on treatment and on outcomes. Note that $0.490.7^{2}=$ 0.49 is around the middle of the distribution of values in table 2, and hence might be considered a moderate or slightly conservative value of the magnitude of control endogeneity. If we impose exogenous controls ($\bar{c}=0$) then we can allow the impact of the omitted variable on outcomes to be 200% as large as the observables and the impact of the omitted variable on treatment to be up to about 240% as large as the observables, and yet still conclude that $\beta_{\text{long}}>0$.
Figure 2: Sensitivity Analysis for Cut Spending on Poor. See body text for discussion.
@@ 6 Conclusion
This paper develops a new approach to assess the sensitivity of regression results to omitted variables that does not require assuming the controls are exogenous. Yet, our approach maintains a key aspect of a popular approach by Altonji et al. (2005) and Oster (2019) which lets researchers calibrate the magnitude of the sensitivity parameter by comparing a measure of selection on observables with a measure of selection on unobservables. Our results are also simple to implement in practice, via the accompanying Stata package regsensitivity. Finally, in our empirical application to Bazzi et al.’s (2020) study of the impact of frontier life on modern culture, we showed that allowing for endogenous controls does matter in practice, leading to richer empirical conclusions than those obtained in Bazzi et al. (2020).

### v6

@@ 5.3 Assessing Selection on Observables
Sensitivity Analysis Based on Oster (2019)
Bazzi et al. (2020) use Oster’s (2019) to assess the impact of omitted variables on their results. That method uses a sensitivity parameter $\delta_{\text{resid}}$, which is informally described in Oster’s appendix A (see appendix C of our companion paper Diegert et al. (2025) for a detailed, formal description of this parameter). The variables $(X,W_{1},W_{2})$ are replaced by $(X^{\perp W_{0}},W_{1}^{\perp W_{0}},W_{2}^{\perp W_{0}})$ to adjust for the non-calibration covariates $W_{0}$. Finally, the auxiliary sensitivity parameter in Oster’s analysis, $R_{Y\sim X,W_{1},W_{2}\mathrel{{\bullet}}W_{0}}^{2}$, is set to $1.3\cdot\widehat{R}_{Y\sim X,W_{1}\mathrel{{\bullet}}W_{0}}^{2}$. Using these parameters and Oster’s analysis, Bazzi et al. (2020) report the estimated breakdown point, which is a measure of how large $\delta_{\text{resid}}$ must be for the estimated effect to be fully explained away by the unobservables.
The second row of Panel B of Table 5.2 shows sample analog estimates of this breakdown point, which is commonly referred to as Oster’s delta. The first row of Panel B shows the values of Oster’s delta as reported in table 2 of Bazzi et al. (2020). These were computed incorrectly. [^5] There are three differences between the two rows in Panel B. First, it appears to us that, rather than using the correct expression in Proposition 3 of Oster (2019), Bazzi et al. (2020) set the first displayed equation on page 193 of Oster (2019) equal to zero and solved for $\delta_{\text{resid}}$. That does not give the correct breakdown point. This mistake was not unique to those authors; while replicating other results, we discovered the same mistake in several other papers published in ‘top 5’ economics journals. Second, Bazzi et al. (2020) used $\widehat{R}_{Y\sim X,W_{0},W_{1}}^{2}$ and $\widehat{R}_{Y\sim X,W_{0}}^{2}$ to adjust for $W_{0}$ rather than $\widehat{R}_{Y\sim X,W_{1}\mathrel{{\bullet}}W_{0}}^{2}$ and $\widehat{R}_{Y\sim X\mathrel{{\bullet}}W_{0}}^{2}$. Third, as we discussed earlier, they include the individual demographics and survey wave fixed effects in $W_{1}$ rather than $W_{0}$. This is a valid choice, but we chose to instead include these variables in $W_{0}$ to keep the covariates in $W_{1}$ constant across columns (1)–(5) for comparability. Based on these estimates, Bazzi et al. (2020) conclude:
@@
The estimates in Table 2 show a substantial range of correlation between the observed covariates in $W_{1}$. Recall that the exogenous controls assumption says that each element of $W_{1}$ is uncorrelated with $W_{2}$, after partialling out $W_{0}$. So under that assumption, if $W_{2}$ was included in this table, it would have a value of zero. Therefore, if $W_{2}$ is a variable similar to the components of $W_{1}$ then we would expect exogenous controls to fail. This suggests that, in this application, robustness conclusions based on $\delta_{\text{resid}}$ could be incorrect. Consequently, one should instead use sensitivity analysis methods which are valid even when the controls are endogenous.
@@
Results from Our Sensitivity Analysis
@@
Outcome: Prefers Cut Public Spending on Poor Republican Presidential Vote Share
@@
Calibrating $\bar{r}_{X}$. Table 5.3 shows estimates of the calibration parameters $\rho_{k}$ that we discussed in section 3.4. Although the treatment variable is the same across all specifications in Table 5.2, the estimation datasets vary. Consequently, the estimates of $\rho_{k}$ vary across specifications as well. Here we only show the estimates corresponding to the specifications in columns (1) and (5) for brevity. First consider column (5), the results for average Republican vote share. Here we see that $\widehat{\rho}_{k}$ is smaller than the $\bar{r}_{X}^{\text{bp}}$ breakdown point estimate of 80.4% for all but one covariate value. This suggests that the results for average Republican vote share is quite robust to selection on unobservables, assuming we view the calibration parameter estimates as plausible values of the unknown ratio $r_{X}$. In contrast, consider column (1), the results for the cut spending on the poor outcome variable. In this case, the $\bar{r}_{X}^{\text{bp}}$ breakdown point estimate of 2.81% is smaller than all the calibration parameters $\widehat{\rho}_{k}$. Even if we use the less conservative breakdown point $\widehat{\bar{r}}^{\text{bp}}=54.1$%, this value is still smaller than half of the calibration parameters. This suggests that the results for cut spending on the poor are sensitive to selection on unobservables, again assuming the calibration parameter estimates are plausible values for the unknown $r_{X}$. Overall, our calibration analysis confirms the findings discussed above: The conclusions using the behavior based outcomes are substantially more robust than the conclusions using questionnaire based outcomes.
Figure 1: Sensitivity Analysis for Average Republican Vote Share. Left: Bounds on $\beta_{\text{long}}$ as a function of $\bar{r}_{X}$. The solid line sets $\bar{r}_{Y}=\infty$ while the dashed line restricts $\bar{r}_{Y}=\bar{r}_{X}$. The tick marks on the horizontal axis at zero show the estimated calibration parameters from Table 5.3. Right: Thick solid line shows an estimate of the breakdown frontier $\bar{r}_{Y}^{\text{bf}}(\bar{r}_{X},0)$, which allows for arbitrarily endogenous controls. The lighter lines show estimated breakdown frontiers for $\underline{c}=0$ and various values of $\overline{c}<1$.
@@
Next we consider robustness to choices of $\bar{r}_{Y}\neq\bar{r}_{X}$. To show this, we plot the estimated breakdown frontier $\bar{r}_{Y}^{\text{bf}}(\bar{r}_{X},0)$ that we showed how to compute in Theorem 2. This is shown as the solid black line in the right plot of figure 1. For any value of $(\bar{r}_{X},\bar{r}_{Y})$ below this line we can conclude that $\beta_{\text{long}}>0$. For example, $(\bar{r}_{X},\bar{r}_{Y})=(2,0.5)$ lies below this line. This means that our conclusion that $\beta_{\text{long}}>0$ is robust to unobservables up to twice as strong as the observables for their impact on treatment, and up to half as strong as the observables for their impact on outcomes.
In section 3.5 we also extended our results to allow researchers to restrict the magnitude of control endogeneity as well. These restrictions shift the breakdown frontier away from zero, as shown by the gray lines in the right plot of figure 1. These plots correspond to the restriction $R_{W_{2}\sim W_{1}\mathrel{{\bullet}}W_{0}}\in[0,\bar{c}]$ for the various displayed values of $\bar{c}$. For example, if we set $\bar{c}=0.7$ then the point $\bar{r}_{X}=\bar{r}_{Y}=110\%$ is now below the breakdown frontier. Hence with this restriction on control endogeneity, our conclusion $\beta_{\text{long}}>0$ is robust to unobservables up to 1.1 times as strong as unobservables in for their impact on treatment and on outcomes. Note that $0.7^{2}=0.49$ is around the middle of the distribution of values in Table 2, and hence might be considered a moderate or slightly conservative value of the magnitude of control endogeneity. If we impose exogenous controls ($\bar{c}=0$) then we can allow the impact of the omitted variable on outcomes to be 200% as large as the observables and the impact of the omitted variable on treatment to be up to about 240% as large as the observables, and yet still conclude that $\beta_{\text{long}}>0$.
Figure 2: Sensitivity Analysis for Cut Spending on Poor. Left: Bounds on $\beta_{\text{long}}$ as a function of $\bar{r}_{X}$. The solid line sets $\bar{r}_{Y}=\infty$ while the dashed line restricts $\bar{r}_{Y}=\bar{r}_{X}$. The tick marks on the horizontal axis at zero show the estimated calibration parameters from Table 5.3. Right: Thick solid line shows an estimate of the breakdown frontier $\bar{r}_{Y}^{\text{bf}}(\bar{r}_{X},0)$, which allows for arbitrarily endogenous controls. The lighter lines show estimated breakdown frontiers for $\underline{c}=0$ and various values of $\overline{c}<1$.
@@ 6 Conclusion
This paper develops a new method to assess the sensitivity of regression results to omitted variables that does not require assuming the controls are exogenous. Our approach maintains a key aspect of the popular methods of Altonji et al. (2005) and Oster (2019) which let researchers calibrate the magnitude of the sensitivity parameter by comparing a measure of selection on observables with a measure of selection on unobservables, while simultaneously resolving several limitations of those earlier methods. Our results are also simple to implement in practice, via the accompanying Stata package regsensitivity. Finally, in our empirical application to Bazzi et al.’s (2020) study of the impact of frontier life on modern culture, we showed that allowing for endogenous controls does matter in practice, leading to richer empirical conclusions than those obtained in Bazzi et al. (2020).
Appendix A Sketch of Main Identification Proof
To provide intuition for our first main identification result, Theorem 1, we give a brief sketch derivation of its proof here. Suppose for simplicity that $W_{2}$ is a scalar. Using properties of correlations and linear projections, it can be shown that the omitted variable bias can be written as
$|\beta_{\text{med}}-\beta_{\text{long}}|=\sqrt{\frac{\operatorname*{var}(Y^{\perp X,W_{1}})}{\operatorname*{var}(X^{\perp W_{1}})}\cdot\frac{R^{2}_{X\sim W_{2}\mathrel{{\bullet}}W_{1}}}{1-R^{2}_{X\sim W_{2}\mathrel{{\bullet}}W_{1}}}\cdot R^{2}_{Y\sim W_{2}\mathrel{{\bullet}}X,W_{1}}}.$ (9)
Versions of this decomposition appear in Hosman et al. (2010) and Cinelli and Hazlett (2020), for example. This expression depends on two unknown partial R-squareds: $R^{2}_{X\sim W_{2}\mathrel{{\bullet}}W_{1}}$ and $R^{2}_{Y\sim W_{2}\mathrel{{\bullet}}X,W_{1}}$. Now suppose we impose Assumption A3, $r_{X}\leq\bar{r}_{X}$. We can write
$r_{X}^{2}=\frac{R^{2}_{X\sim W_{1},W_{2}}-R^{2}_{X\sim W_{1}}}{R^{2}_{X\sim W_{1},W_{2}}-R^{2}_{X\sim W_{2}}}=\frac{R_{X\sim W_{2}\mathrel{{\bullet}}W_{1}}^{2}(1-R^{2}_{X\sim W_{1}})}{R_{X\sim W_{2}\mathrel{{\bullet}}W_{1}}^{2}(1-R^{2}_{X\sim W_{1}})+R^{2}_{X\sim W_{1}}-R^{2}_{X\sim W_{2}}}$ (10)
and hence the assumption that $r_{X}\leq\bar{r}_{X}$ does restrict the value of $R_{X\sim W_{2}\mathrel{{\bullet}}W_{1}}^{2}$. For example, setting $\bar{r}_{X}=0$ implies $R^{2}_{X\sim W_{2}\mathrel{{\bullet}}W_{1}}=0$. Equation (10) does not depend on $R^{2}_{Y\sim W_{2}\mathrel{{\bullet}}X,W_{1}}$, however. Since equation (10) is monotonically increasing in $R^{2}_{Y\sim W_{2}\mathrel{{\bullet}}X,W_{1}}$, we can therefore set it to 1 to obtain the worst case bounds (this can be done independently from $R^{2}_{X\sim W_{2}\mathrel{{\bullet}}W_{1}}$ by Lemma LABEL:lem:unconstrainedR2_vector). Similarly, equation (10) is monotonically increasing in $R_{X\sim W_{2}\mathrel{{\bullet}}W_{1}}^{2}$. So it suffices to find the largest possible value of this R-squared that is consistent with the assumption that $r_{X}\leq\bar{r}_{X}$. To do this, notice that equation (10) also depends on the unknown $R^{2}_{X\sim W_{2}}$. This term does not enter the OVB equation (9) and hence we can treat it as a free parameter and simply pick the largest value of $R^{2}_{X\sim W_{2}\mathrel{{\bullet}}W_{1}}\in[0,1)$ consistent with some value of $R^{2}_{X\sim W_{2}}\in[0,1]$ and with the observed data and the constraint that equation (10) is not larger than $\bar{r}_{X}$. The solution is $R^{2}_{X\sim W_{2}}=0$ and
$R^{2}_{X\sim W_{2}\mathrel{{\bullet}}W_{1}}$ $=\begin{cases}\frac{\bar{r}_{X}^{2}R^{2}_{X\sim W_{1}}}{(1-R^{2}_{X\sim W_{1}})(1-\bar{r}_{X}^{2})}&\text{ if }\bar{r}_{X}^{2}<1-R^{2}_{X\sim W_{1}}\\ 1&\text{ if }\bar{r}_{X}^{2}\geq 1-R^{2}_{X\sim W_{1}}\end{cases}$
$=\min\left\{\frac{\bar{r}_{X}^{2}R^{2}_{X\sim W_{1}}}{(1-R^{2}_{X\sim W_{1}})(1-\bar{r}_{X}^{2})},1\right\}.$
See Lemmas LABEL:lem:unconstrainedR2_vector, LABEL:lem:marg_R2_XonW2_vector, and LABEL:lem:max_abstract_vector for formal details. Substituting this and $R^{2}_{Y\sim W_{2}\mathrel{{\bullet}}X,W_{1}}=1$ into (9) gives
$|\beta_{\text{med}}-\beta_{\text{long}}|$ $=\sqrt{\dfrac{\operatorname*{var}(Y^{\perp X,W_{1}})}{\operatorname*{var}(X^{\perp W_{1}})}\cdot\frac{\min\left\{\frac{\bar{r}_{X}^{2}R^{2}_{X\sim W_{1}}}{(1-R^{2}_{X\sim W_{1}})(1-\bar{r}_{X}^{2})},1\right\}}{1-\min\left\{\frac{\bar{r}_{X}^{2}R^{2}_{X\sim W_{1}}}{(1-R^{2}_{X\sim W_{1}})(1-\bar{r}_{X}^{2})},1\right\}}\cdot 1}$
$=\begin{cases}\sqrt{\dfrac{\operatorname*{var}(Y^{\perp X,W_{1}})}{\operatorname*{var}(X^{\perp W_{1}})}\cdot\dfrac{\bar{r}_{X}^{2}R^{2}_{X\sim W_{1}}}{1-R^{2}_{X\sim W_{1}}-\bar{r}_{X}^{2}}}&\text{ if }\bar{r}_{X}^{2}<1-R^{2}_{X\sim W_{1}}\\ +\infty&\text{ if }\bar{r}_{X}^{2}\geq 1-R^{2}_{X\sim W_{1}}.\end{cases}$
This is the largest magnitude of omitted variable bias under the restriction that $r_{X}\leq\bar{r}_{X}$, which therefore gives us the bounds in Theorem 1. This argument can be modified to account for constraints on $r_{Y}$ as in Theorem 2 or constraints on the correlation between $W_{1}$ and $W_{2}$ as in Theorems 3 and 4. We omit this discussion for brevity; see the online appendices for full proofs.
Appendix B Additional Empirical Analysis
The Effect of the Choice of Calibration Covariates
Figure 3: Effect of Calibration Covariates on Analysis For Republican Vote Share. See body text for discussion.
In section 3.4 we discussed the importance of choosing which variables to calibrate against (the variables in $W_{1}$) versus which variables to use as controls only (the variables in $W_{0}$). Here we briefly illustrate this in our empirical application. The results in Table 5.2 and figures 1 and 2 all include state fixed effects as controls, but do not use them for calibration; that is, these variables are in $W_{0}$. Next we consider the impact of instead putting them into $W_{1}$ and calibrating the magnitude of selection on unobservables against them, in addition to the geographic and climate controls already in $W_{1}$.
Figure 3 shows plots corresponding to figure 1, but now also using state fixed effects for calibration. We first see that the identified sets for $\beta_{\text{long}}$ (left plot) are larger, for any fixed $\bar{r}_{X}$. This makes sense because the interpretation of $\bar{r}_{X}$ has changed with the change in calibration controls. In particular, the breakdown point $\bar{r}_{X}^{\text{bp}}$ is now about 30%, whereas previously it was about 80%. By including state fixed effects—which have a large amount of explanatory power—in our calibration controls, we have increased the magnitude of selection on observables. Holding selection on unobservables fixed, this implies that $r$ must decrease. This discussion reiterates the point that the magnitude of $\bar{r}_{X}$ must always be interpreted as dependent on the set of calibration controls. For example, our finding in figure 3 that the estimated $\bar{r}_{X}^{\text{bp}}$ is about 30% should not be interpreted as saying that the results are sensitive; in fact, an effect about 30% as large as these calibration covariates is substantially large, and so it may be that we do not expect the omitted variable to have such a large additional impact.
The right plot in figure 3 shows the estimated breakdown frontiers. The frontiers have all shifted inward, compared to the right plot of figure 1 which did not use state fixed effects for calibration. Consequently, a superficial reading of this plot may suggest that the results for average Republican vote share are no longer robust. However, as we just emphasized in our discussion of the left plot, by including state fixed effects in the calibration covariates $W_{1}$, we are changing the meaning of all sensitivity parameters. Since the expanded set of calibration covariates has substantial explanatory power, even a relaxation like $(\bar{r}_{Y},\bar{r}_{X})=(50\%,50\%)$—which is below the breakdown frontier and hence allows us to conclude that $\beta_{\text{long}}$ is positive—could be considered to be a large impact of omitted variables. So these figures do not change our overall conclusions about the robustness of the analysis for average Republican vote share.
Appendix C The Identified Set For $\beta_{\text{long}}$ With Fixed $(\bar{r}_{X},\bar{r}_{Y},\underline{c},\bar{c}$)
In this appendix we characterize the identified set for $\beta_{\text{long}}$, the coefficient on $X$ in the long regression of $Y$ on $(1,X,W_{1},W_{2})$, under assumptions A3A6 hold and when $W_{2}$ is a scalar. That is, we use information from all three sensitivity parameters to learn about $\beta_{\text{long}}$ (Theorem 5 below). This result is a preliminary step in deriving our main results in sections 3.2 and 3.3.
Here and throughout the appendix we let $\|a\|_{V}\coloneqq\sqrt{a^{\prime}Va}$ denote the weighted Euclidean norm, for a positive definite matrix $V$. Let $\|\cdot\|$ denote the unweighted Euclidean norm. Let $\Sigma_{\text{obs}}\coloneqq\operatorname*{var}(W_{1})$. Define
[…]
where these equations are
$\operatorname*{cov}(Y,X)$ $=b\operatorname*{var}(X)+g_{1}^{\prime}(\operatorname*{var}(W_{1})+c\tilde{r}_{X}^{\prime}+\tilde{r}_{Y}c^{\prime}+\tilde{r}_{Y}\tilde{r}_{X}^{\prime})p_{1}$ (12)
$\operatorname*{cov}(Y,W_{1})$ $=b\operatorname*{cov}(X,W_{1})+g_{1}^{\prime}(\operatorname*{var}(W_{1})+\tilde{r}_{Y}c^{\prime})$ (13)
$\operatorname*{cov}(X,W_{1})$ $=p_{1}^{\prime}(\operatorname*{var}(W_{1})+\tilde{r}_{X}c^{\prime})$ (14)
$\operatorname*{var}(Y)$ $>b^{2}\operatorname*{var}(X)+g_{1}^{\prime}\left(\operatorname*{var}(W_{1})+\tilde{r}_{Y}\tilde{r}_{Y}^{\prime}+2\tilde{r}_{Y}c^{\prime}\right)g_{1}$
$\quad+2bg_{1}^{\prime}(\operatorname*{var}(W_{1})+c\tilde{r}_{X}^{\prime}+\tilde{r}_{Y}c^{\prime}+\tilde{r}_{Y}\tilde{r}_{X}^{\prime})p_{1}$ (15)
$\operatorname*{var}(X)$ $>p_{1}^{\prime}(\operatorname*{var}(W_{1})+2\tilde{r}_{X}c^{\prime}+\tilde{r}_{X}\tilde{r}_{X}^{\prime})p_{1}$ (16)
$1$ $>c^{\prime}\operatorname*{var}(W_{1})^{-1}c.$ (17)
Theorem 5.
Suppose the joint distribution of $(Y,X,W_{1})$ is known. Suppose A1 holds. Suppose A3A6 hold. Suppose $W_{2}$ is a scalar with $\operatorname*{var}(W_{2})$ normalized to 1. Let $\mathcal{B}_{I}(\bar{r}_{X},\bar{r}_{Y},\underline{c},\bar{c})$ denote the identified set for $\beta_{\text{long}}$ under these assumptions. Then
$\mathcal{B}_{I}(\bar{r}_{X},\bar{r}_{Y},\underline{c},\bar{c})=\bigcup_{(\tilde{r}_{X},\tilde{r}_{Y},c):\|\tilde{r}_{X}\|_{\Sigma_{\text{obs}}^{-1}}\leq\bar{r}_{X},\|\tilde{r}_{Y}\|_{\Sigma_{\text{obs}}^{-1}}\leq\bar{r}_{Y},\|c\|_{\Sigma_{\text{obs}}^{-1}}\in[\underline{c},\bar{c}]}\mathcal{B}(\tilde{r}_{X},\tilde{r}_{Y},c).$ (18)